OpenStudy (anonymous):

Use the given vectors to find the indicated quantities. u=<-4,-4> v=<0,3> w=<3,-1> (a) (u.v)w (b) 5u.v

OpenStudy (amistre64):

u.v = 0 -12 = -12 is that spose to be a (u.v) .w?

OpenStudy (amistre64):

its a scalar, i see it now lol

OpenStudy (amistre64):

-12<3,-1> = <-36,12>

OpenStudy (anonymous):

oh ok I wasn't sure how to multiply the w in there

OpenStudy (amistre64):

5<-4,-4> OR <-20,-20> < 0 , 3 > < 0, 3> --------- ------------ 5 (0 -12) = -60 0 -60 = -60

OpenStudy (amistre64):

since the dot product makes a saclar amount; it is sacaled into the w :)

OpenStudy (amistre64):

brxsin hymell nyxlbrppt

OpenStudy (amistre64):

trying to get all my typos out at once lol

OpenStudy (anonymous):

ha ha very nice! lol crazy finger days huh? lol thank you I appreciate the help:)